RC meeting and Run Plan, Tuesday Sept. 19, 2023

From SBS wiki
Jump to: navigation, search

Return to the RC Meeting Calendar

RC Meeting Information

Time:at 3:30pm today

  • Location: Counting house conference room
  • ZoomGov Meeting Connection Information


Useful Links


Hall A call-in instructions:

  • Call RC first.
  • Techs on site 7-4:30 Mon-Th, 7-11AM Fr. RC will call Hall A work coordinator (Zak) during normal working hours. After hours, RC calls Tech on-call.
  • After 9pm everything non-emergency needs to be deferred. What is emergency and what's not? Call RC. RC talks with Mark to decide.

Hall Access procedure:

  • Access check list (binder on the counting room desk) needs to be filled out before making an access.
  • Scan (take pictures with phone) and post it to the logbook.
  • Any work inside the SBS, BB boundary need to have PS locked out -- requires two QEWs to throw the big switch. Exception: field measurements.

New instruction for shift crew when changing target:

  • Beam restoration instructions have been updated. Printed version is in the counting room.
  • When we chnage target, make sure the SOFs of Target BPMs are set to the correct values.
  • SOFs for different targets: https://logbooks.jlab.org/entry/4179164 (Note: SOFs for "No Target" has been updated and they are same as the He3 target values)

Readiness for Beam

Acclerator status

  • A few hours of downtime during swing: 1C05 trip magnet regulation fault, PS replaced (https://logbooks.jlab.org/entry/4182290). ABU+BANU: 2.5+1.0 = 3.5hrs.
  • Downtime during the daytime mostly due to RF issues. AB+BANU: 5.1+0.6 = 5.7hrs
  • Took 30uA beam for swing and owl. Got 30-36uA during the day shift. ABU+BANU: 3.1+0.7 = 3.8hrs
  • Charge collected:

- Swing: 0.225C - Owl: 0.288C - Day: 0.244C

Subsystems Checklist

For each subsystem we ask that the responsible person go over the following points:

  1. The overall status of your subsystem
  2. Will your subsystem be ready well before Hall lock-up on Friday or at the latest by Monday morning?
  3. Does your system being ready depend on your team only or are outside resources required?


  • SBS magnet and Bigbite magnets
    • work as expected
  • Corrector magnets:
    • work as expected
  • Target status (magnets, measurements, laser, installation, alignment, polarization, power etc)
  • HCal (Jiwan Poudel)
    • VME DAC GUI does not read the correct threshold -- pending issue to be fixed
  • Bigbite (Arun Tadepalli)
  • Counting House Alarms (Bill Henry)
  • DAQ (Alex Camsonne)
  • GEMs (Holly Szumila-Vance)
  • Moller polarimeter (Don Jones)
    • Hall C is going to do the Moller measurement tomorrow.
  • Hodoscopes (Rachel Montgomery):
    • Hodo status ok. Will fix two bottom bars with low efficiency.

Task list during the access, Tuesday 9/20

  • Please let SL know that you are going down to the hall, where to work, and after you are done.
  • Make a HALOG entry summarizing the work
  • Start early in the morning if possible, to reduce the interference with the laser work (put the hall in CANS mode)
  • List:
    • SBS PS acid flushing. Crew will be there as soon as we can get into the hall
    • Target work
    • Sean: BB GEM Layer 3 (~2h)
    • Jiwan: HCal (a few channels missing TDC)
    • Alex: TS crate, GRINCH FAN VXS reboot, threshold RPi monitor
    • Ellen: Grinch bottle weight readback
    • Ralph: Hodoscope (bottom two blocks with low efficiency)
    • Don: Moller target magnet
    • Jimmy: ECal HV crates. DSG will do some work on the ECal heater prototype

Shift schedule

  • Looks much better! Just a couple of slots to fill next week (DAY, SWING).

Run plan

  • Take production data until tomorrow morning, up to 36uA.
  • At around 6:50am tomorrow, beam off and ramp down SBS and BB magnets before the access.
  • Hall open around 7:30-8AM.
  • After the access, restore the CW beam. Check Carbon hole data and do the target centering.
  • Also check the detector plots in the meantime.
  • Bring up the current up to 45uA on He3 if possible.
  • Take data with glass cells overnight.

List of items in case target isn’t ready:

  1. Take data on reference cell at 10 atm H2 (portions that add up to 1 day)
  2. BigBite Sieve slit => only during day (Techs needed)
    1. with 8 carbon foils and take optics for 1 hr at 40 uA (Sean/Andrew - confirm time)
    2. 1 hour for single foil (magnets off) => Chance BBcal HV setting to appropriate values
  3. Moller measurement
  4. BigBite/single foil location reconstruction (~20 mins?) - BigBite + HCAL detector experts on call consulting analysis experts like Andrew/Sean (magnet on)